# Roles of the scalar $f_{0}(500)$ and $f_{0}(980)$ in the process $D^{0}\to\pi^{0}\pi^{0}\bar{K}^{0}$

### v1

@@ Abstract
Motivated by the near-threshold enhancement and the dip structure around 1 GeV in the $\pi^{0}\pi^{0}$ invariant mass distribution of the process $D^{0}\to\pi^{0}\pi^{0}\bar{K}^{0}$ observed by the CLEO Collaboration, we have investigated this process by taking into account the contribution from the $S$-wave pseudoscalar meson-pseudoscalar meson interactions within the chiral unitary approach, and also the one from the intermediate resonance $K^{*}(892)$. Our results are in good agreement with the CLEO measurements, which implies that, the near-threshold enhancement near the $\pi^{0}\pi^{0}$ threshold is mainly due to the contributions from the scalar meson $f_{0}(500)$ and the intermediate $K^{*}$, and the cusp structure around 1 GeV in the $\pi^{0}\pi^{0}$ invariant mass distribution should be associated with the scalar meson $f_{0}(980)$.
@@ I Introduction
The hadronic decays of the charmed hadrons could provide valuable insights for understanding the short distance weak interaction and the long-distance Quantum Chromodynamics (QCD) interaction Wang:2021ail ; Lyu:2024qgc ; Li:2024rqb ; Zhang:2024jby . Especially, the final state interaction are the key ingredients in the production of the light scalar mesons, i.e. $f_{0}(500)$, $f_{0}(980)$, and $a_{0}(980)$, which are of particular interest for exploring the nature of the light scalar mesons Liang:2016hmr ; Debastiani:2016ayp ; Oset:2016lyh ; Wang:2021naf ; Wang:2022nac ; Feng:2020jvp ; Wang:2020pem .
@@
Recently, BESIII, BaBar, and Belle/Belle II have accumulated a lot of experimental data about the multi-body decays of the charmed hadrons, which provides an important platform to investigate these light scalar mesons Wang:2021kka ; Liang:2014ama ; Xie:2014gla ; BaBar:2010nhz ; Ling:2021qzl ; Ding:2024lqk ; Ding:2023eps . For instance, the BESIII and BABAR Collaborations have performed the amplitude analysis of the process $D_{s}^{+}\to K^{+}K^{-}\pi^{+}$, and observed the near-threshold enhancement in the $K^{+}K^{-}$ invariant mass distribution BESIII:2020ctr ; BaBar:2010wqe . In Ref. Wang:2021naf , the authors have investigated this process by taking into account the $S$-wave pseudoscalar-pseudoscalar interactions within the unitary chiral approach, and concluded that scalar meson $f_{0}(980)$ gives the dominant contribution for the near-threshold enhancement in the $K^{+}K^{-}$ invariant mass distribution. In addition, the authors of Ref. Zhang:2022xpf have studied the decay $D^{+}_{s}\to\pi^{+}\pi^{0}\pi^{0}$, and concluded that the cusp structure around 1 GeV in the $\pi^{0}\pi^{0}$ invariant mass distribution could be associated with the scalar meson $f_{0}(980)$.
@@
In 2011, the CLEO Collaboration has conducted a dalitz plot analysis on $D^{0}\to K_{S}^{0}\pi^{0}\pi^{0}$ using data set of 818 pb-1 of $e^{+}e^{-}$ collisions accumulated at $\sqrt{s}=3.77$ GeV CLEO:2011cnt , and concluded that combined $\pi^{0}\pi^{0}$ S-wave contributes $(28.9\pm 6.3\pm 3.1)\%$ of the total decay rate, while the $D^{0}\to\bar{K}^{*0}\pi^{0}$ contributes $(65.6\pm 5.3\pm 2.5)\%$. In addition, the decay branching fraction of the $D^{0}\to K^{0}\pi^{0}\pi^{0}$ was also measured to be $(1.058\pm 0.038\pm 0.073)\%$ CLEO:2011cnt . In the $\pi^{0}\pi^{0}$ invariant mass distribution, a cusp structure followed by a dip appears around 1 GeV, which may be associated with the scalar meson $f_{0}(980)$, and an enhancement was also found near the $\pi^{0}\pi^{0}$ threshold. Meanwhile, one could find a clear peak of $K^{*}(892)$ in the $\pi^{0}K$ invariant mass distribution. Due to the conservation of the charge parity, there is no contribution from the intermediate vector meson $\rho$ in the process $D^{0}\to\bar{K}^{0}\pi^{0}\pi^{0}$, thus this mode could provide a more clear way to investigate the $S$-wave $\pi\pi$ substructure than the process $D^{0}\to K^{0}_{s}\pi^{+}\pi^{-}$ BaBar:2005rek ; CLEO:2002uvu .
@@
In this work, we will investigate the process $D^{0}\to\pi^{0}\pi^{0}\bar{K}^{0}$ by considering the scalar $f_{0}(980)$ generated from the $S$-wave pseudoscalar-pseudoscalar meson interaction within the chiral unitary approach. In addition, we also consider the contribution from the intermediate vector meson $K^{*}(892)$.
@@
In Sec. II.1, we first present the mechanism of the Cabibbo-favored process $D^{0}\to\bar{K}^{0}\pi^{0}\pi^{0}$ via the $S$-wave pseudoscalar meson-pseudoscalar meson interactions, which could dynamically generated the scalar mesons $f_{0}(500)$ and $f_{0}(980)$. Meanwhile, this process could also happen via the intermediate vector meson $K^{*}(892)$, and the corresponding formalism and the double differential decay width are given in Sec. II.2.
@@ II.1 $S$-wave pseudoscalar-pseudoscalar interactions
Figure 1: Feynman diagram of $D^{0}\to K^{0}u\bar{u}$ and hadronization of the $u\bar{u}$ through $q\bar{q}$ with vacuum quantum numbers.
@@
where we adopt the approximate $\eta-\eta^{\prime}$ mixing Lyu:2023ppb ; Bramon:1992kr . Then all the possible meson-meson components of Eq. (1) are,
@@
with the pion decay constant $f=93$ MeV and the invariant mass square $s$ of the meson-meson system. The masses of the mesons are taken from the Review of Particle Physics (RPP) ParticleDataGroup:2022pth . We have take the unitary normalization $\left|\eta\eta\right>\to\dfrac{1}{\sqrt{2}}\left|\eta\eta\right>$, $\left|\pi^{0}\pi^{0}\right>\to\dfrac{1}{\sqrt{2}}\left|\pi^{0}\pi^{0}\right>$ to account for the identity of particles when using the $G$ function without an extra factor in Eq. (7Liang:2014tia . The loop function for two-meson system is,
@@
where the $m_{1}$, $m_{2}$ are the masses of mesons in the $l$-channel, $q$ is the four-momentum of one meson, and the $P$ is the total four-momentum, with $s=P^{2}$. Since the loop function $G$ in Eq. (9) is logarithmically divergent Oller:1998hw , we have two methods to solve this singular integral, either using the three-momentum cut-off method, or the dimensional regularization method. In this work, the integral of the loop function is performed integrating exactly for $q^{0}$ and implementing a cut off $q_{\rm max}$ for the three-momentum Oller:1997ti ,
@@
In addition, the most reliable prediction range of the $S$-wave pseudoscalar meson-pseudoscalar meson interaction within the chiral unitary approach is up to $1100\sim 1200$ MeV, and one can not use this model for higher invariant masses. In order to present a reasonable $\pi^{0}\pi^{0}$ invariant mass distribution, we evaluate the $Gt$ combinations of Eq. (5) up to $M_{\rm cut}=1.1$ GeV, and from there we multiply $Gt$ by a smooth factor to make it gradually decreasing at large $M_{\pi^{0}\pi^{0}}$, as done in Ref. Debastiani:2016ayp ,
@@ III Results and discussions
First, we have performed a fit to the experimental data of the $\pi^{0}\pi^{0}$ and $\pi^{0}K^{0}$ invariant mass distributions CLEO:2011cnt . Then we obtained the $\chi^{2}/d.o.f.$ is 237.8/(47+48-2)=2.56, and the fitted parameters $(V^{\prime}_{p})^{2}=(6.275\pm 0.2472)\times 10^{9}$, $\beta=(9.5195\pm 0.313)\times 10^{-2}$.
Next, we have calculated the $\pi^{0}\pi^{0}$ and $\pi^{0}\bar{K}^{0}$ invariant mass distribution with the fitted parameters in Fig. 4. The black points with error bars labeled as ‘CLEO data’ are the CLEO-c data taken from Ref. CLEO:2011cnt . One can find that, in the $\pi^{0}\pi^{0}$ invariant mass distribution, our results could well reproduce the structure of the cusp followed by a dip, which could be associated with the scalar meson $f_{0}(980)$, dynamically generated from the $S$-wave pseudoscalar meson pseudoscalar meson interactions. Meanwhile, our results show that the peak around $M^{2}_{\pi^{0}\pi^{0}}=0.2$ GeV2 is mainly due to the scalar meson $f_{0}(500)$ and the reflection of the vector meson $\bar{K}^{*0}$.
In addition, our results of the $\pi\bar{K}^{0}$ invariant mass distribution are in good agreement with the CLEO measurements. The narrow peak structure around $M^{2}_{\pi^{0}\bar{K}^{0}}=0.7$ GeV2 should be associated with the vector meson $K^{*}{(892)}$, while the contribution from the $S$-wave pseudoscalar meson-pseudoscalar meson interaction serves as the smooth background in the $\pi\bar{K}^{0}$ invariant mass distribution.
@@
We also present the Dalitz plots of ‘${M^{2}_{\pi^{0}\bar{K}^{0}}}$’ and ‘${M^{2}_{\pi^{0}\pi^{0}}}$’ in Fig. 5, which clearly shows the structures of $f_{0}(980)$ and $K^{*}(892)$. One can also find that the peak near the $\pi^{0}\pi^{0}$ threshold is mainly due to the scalar meson $f_{0}(500)$ and the reflection of the vector meson $\bar{K}^{*0}$.
Finally, with the fitted parameters, we could estimate that the $S$-wave $\pi^{0}\pi^{0}$ interaction contributes about $47.7\%$ of the total decay rate, larger than the experimental data $(28.9\pm 6.3\pm 3.1)\%$, while the intermediate resonance $K^{*}(892)$ contributes about $52.3\%$ of the total decay rate, consistent with the experimental data $(65.6\pm 5.3\pm 2.5)\%$ within the uncertainties CLEO:2011cnt . It should be pointed that the contribution from the scalar $f_{0}(500)$ was not taken into account by CLEO CLEO:2011cnt , which plays an important role for the near-threshold enhancement of the $\pi^{0}\pi^{0}$ invariant mass distribution.
Figure 5: The Dalitz plots of ’$M^{2}_{\pi^{0}\pi^{0}}$’ vs.’$M^{2}_{\pi^{0}\bar{K}^{0}}$’ for the process $D^{0}\to\pi^{0}\pi^{0}\bar{K}^{0}$.
@@ IV SUMMARY AND CONCLUSION
Motivated by the CLEO measurements about the process $D^{0}\to\pi^{0}\pi^{0}\bar{K}^{0}$ CLEO:2011cnt , we have investigated this process by taking into account the contribution from the $S$-wave pseudoscalar meson-pseudoscalar meson interaction within the chiral unitary method, which could dynamically generate intermediate resonance states $f_{0}(980)$ and $f_{0}(500)$, and also the contribution from the intermediate vector meson $\bar{K}^{*}{(892)}$.
First, we have performed a fit to the experimental data of the $\pi^{0}\pi^{0}$ and $\pi^{0}K_{S}^{0}$ invariant mass distributions. With the fitted parameters, our results could well reproduce the dip structure followed a dip around 1 GeV in the $\pi^{0}\pi^{0}$ invariant mass distribuiton, which is mainly due to the contribution of scalar $f_{0}(980)$. Our results indicate that the peak near the $\pi^{0}\pi^{0}$ threshold is due to the scalar meson $f_{0}(500)$ and the reflection of the $\bar{K}^{*}$. Our results of the $\pi^{0}\bar{K}^{0}$ invariant mass distribution are in good agreement with the CLEO data, which supports that the peak around 892 MeV in the $\pi^{0}\bar{K}^{0}$ invariant mass distribution should be associated with $K^{*}{(892)}$.

### v2

@@ Abstract
Motivated by the near-threshold enhancement and the dip structure around 1 GeV in the $\pi^{0}\pi^{0}$ invariant mass distribution of the process $D^{0}\to\pi^{0}\pi^{0}\bar{K}^{0}$ observed by the CLEO Collaboration, we have investigated this process by taking into account the contribution from the $S$-wave pseudoscalar meson-pseudoscalar meson interactions within the chiral unitary approach, and also the one from the intermediate resonance $\bar{K}^{*}(892)$. Our results are in good agreement with the CLEO measurements, which implies that, the near-threshold enhancement near the $\pi^{0}\pi^{0}$ threshold is mainly due to the contributions from the scalar meson $f_{0}(500)$ and the intermediate $\bar{K}^{*}$, and the cusp structure around 1 GeV in the $\pi^{0}\pi^{0}$ invariant mass distribution should be associated with the scalar meson $f_{0}(980)$.
@@ I Introduction
The hadronic decays of the charmed hadrons could provide valuable insights for understanding the short distance weak interaction and the long-distance Quantum Chromodynamics (QCD) interaction Wang:2021ail ; Lyu:2024qgc ; Duan:2024led ; Li:2024rqb ; Zhang:2024jby . Especially, the final state interaction are the key ingredients in the production of the light scalar mesons, i.e. $f_{0}(500)$, $f_{0}(980)$, and $a_{0}(980)$, which are of particular interest for exploring the nature of the light scalar mesons Liang:2016hmr ; Debastiani:2016ayp ; Oset:2016lyh ; Wang:2021naf ; Wang:2022nac ; Feng:2020jvp ; Wang:2020pem .
@@
Recently, BESIII, BABAR, and Belle/Belle II have accumulated a lot of experimental data about the multi-body decays of the charmed hadrons, which provides an important platform to investigate these light scalar mesons Wang:2021kka ; Liang:2014ama ; Xie:2014gla ; BaBar:2010nhz ; Ling:2021qzl ; Ding:2024lqk ; Ding:2023eps . For instance, the BESIII and BABAR Collaborations have performed the amplitude analysis of the process $D_{s}^{+}\to K^{+}K^{-}\pi^{+}$, and observed the near-threshold enhancement in the $K^{+}K^{-}$ invariant mass distribution BESIII:2020ctr ; BaBar:2010wqe . In Ref. Wang:2021naf , the authors have investigated this process by taking into account the $S$-wave pseudoscalar-pseudoscalar interactions within the chiral unitary approach, and concluded that scalar meson $f_{0}(980)$ gives the dominant contribution for the near-threshold enhancement in the $K^{+}K^{-}$ invariant mass distribution. In addition, the authors of Ref. Zhang:2022xpf have studied the decay $D^{+}_{s}\to\pi^{+}\pi^{0}\pi^{0}$, and concluded that the cusp structure around 1 GeV in the $\pi^{0}\pi^{0}$ invariant mass distribution could be associated with the scalar meson $f_{0}(980)$, which favors the no-$q\bar{q}$ nature of $f_{0}(980)$.
@@
In 2011, the CLEO Collaboration has conducted a Dalitz plot analysis on $D^{0}\to K_{S}^{0}\pi^{0}\pi^{0}$ using data set of 818 pb-1 of $e^{+}e^{-}$ collisions accumulated at $\sqrt{s}=3.77$ GeV CLEO:2011cnt , and concluded that combined $\pi^{0}\pi^{0}$ $S$-wave contributes $(28.9\pm 6.3\pm 3.1)\%$ of the total decay rate, while the $D^{0}\to\bar{K}^{*0}\pi^{0}$ contributes $(65.6\pm 5.3\pm 2.5)\%$. In addition, the decay branching fraction of the $D^{0}\to\bar{K}^{0}\pi^{0}\pi^{0}$ was also measured to be $(1.058\pm 0.038\pm 0.073)\%$ CLEO:2011cnt . In the $\pi^{0}\pi^{0}$ invariant mass distribution, a cusp structure followed by a dip appears around 1 GeV, which may be associated with the scalar meson $f_{0}(980)$, and an enhancement was also found near the $\pi^{0}\pi^{0}$ threshold. Meanwhile, one could find a clear peak of $\bar{K}^{*}(892)$ in the $\pi^{0}\bar{K}$ invariant mass distribution. Due to the conservation of the charge parity, there is no contribution from the intermediate vector meson $\rho$ in the process $D^{0}\to\bar{K}^{0}\pi^{0}\pi^{0}$, thus this mode could provide a more clear way to investigate the $S$-wave $\pi\pi$ substructure than the process $D^{0}\to K^{0}_{s}\pi^{+}\pi^{-}$ BaBar:2005rek ; CLEO:2002uvu .
@@
In this work, we will investigate the process $D^{0}\to\pi^{0}\pi^{0}\bar{K}^{0}$ by considering the scalar $f_{0}(980)$ generated from the $S$-wave pseudoscalar-pseudoscalar meson interaction within the chiral unitary approach. In addition, we also consider the contribution from the intermediate vector meson $\bar{K}^{*}(892)$.
@@
In Sec. II.1, we first present the mechanism of the Cabibbo-favored process $D^{0}\to\bar{K}^{0}\pi^{0}\pi^{0}$ via the $S$-wave pseudoscalar meson-pseudoscalar meson interactions, which could dynamically generated the scalar mesons $f_{0}(500)$ and $f_{0}(980)$. Meanwhile, this process could also happen via the intermediate vector meson $\bar{K}^{*}(892)$, and the corresponding formalism and the double differential decay width are given in Sec. II.2.
@@ II.1 $S$-wave pseudoscalar-pseudoscalar interactions
where we adopt the approximate $\eta-\eta^{\prime}$ mixing Lyu:2023ppb ; Bramon:1992kr  [^1] According to Review of Particle Physics (PRR) ParticleDataGroup:2022pth , the mixing angle is between $-10^{\circ}$ and $-20^{\circ}$, and a recent Lattice calculations support the value $\theta_{P}=-14.1^{\circ}\pm 2.8^{\circ}$ by reproducing the masses of the $\eta$ and $\eta^{\prime}$ Christ:2010dd . In this work, we adopt the mixing from Ref. Bramon:1992kr , as follows, $\eta\sim\frac{1}{\sqrt{3}}(u\bar{u}+d\bar{d}-s\bar{s}),$ $\eta^{\prime}\sim\frac{1}{\sqrt{6}}(u\bar{u}+d\bar{d}+2s\bar{s}).$. Then all the possible meson-meson components of Eq. (1) are,
@@
Figure 1: Feynman diagram of $D^{0}\to\bar{K}^{0}u\bar{u}$ and hadronization of the $u\bar{u}$ through $q\bar{q}$ with vacuum quantum numbers.
@@
with the pion decay constant $f=93$ MeV and the invariant mass square $s$ of the meson-meson system. The masses of the mesons are taken from RPP ParticleDataGroup:2022pth . We have take the unitary normalization $\left|\eta\eta\right>\to\dfrac{1}{\sqrt{2}}\left|\eta\eta\right>$, $\left|\pi^{0}\pi^{0}\right>\to\dfrac{1}{\sqrt{2}}\left|\pi^{0}\pi^{0}\right>$ to account for the identity of particles when using the $G$ function without an extra factor in Eq. (7Liang:2014tia . The loop function for two-meson system is,
@@
where the $m_{1}$, $m_{2}$ are the masses of mesons in the $l$-channel, $q$ is the four-momentum of one meson, and the $P$ is the total four-momentum, with $s=P^{2}$. Since the loop function $G$ in Eq. (9) is logarithmically divergent Oller:1998hw , we have two methods to solve this singular integral, either using the three-momentum cut-off method, or the dimensional regularization method, by introducing free parameters, which are determined by giving rise to the pole position of the resonance. In this work, the integral of the loop function is performed integrating exactly for $q^{0}$ and implementing a cut off $q_{\rm max}$ for the three-momentum Oller:1997ti ,
@@
In addition, the most reliable prediction range of the $S$-wave pseudoscalar meson-pseudoscalar meson interaction within the chiral unitary approach is up to $1100\sim 1200$ MeV, and one can not use this model for higher invariant masses. In order to present a reasonable $\pi^{0}\pi^{0}$ invariant mass distribution, we evaluate the $Gt$ combinations of Eq. (5) up to $M_{\rm cut}=1.1$ GeV, and from there on we multiply $Gt$ by a smooth factor to make it gradually decreasing at large $M_{\pi^{0}\pi^{0}}$, as done in Ref. Debastiani:2016ayp ,
@@ III Results and discussions
First, we have performed a fit to the experimental data of the $\pi^{0}\pi^{0}$ and $\pi^{0}K^{0}$ invariant mass distributions CLEO:2011cnt . Then we obtain $\chi^{2}/d.o.f.$ is 237.8/(47 + 48-2) = 2.56, and the fit parameters $(V^{\prime}_{p})^{2}=(6.275\pm 0.2472)\times 10^{9}$, $\beta=(9.5195\pm 0.313)\times 10^{-2}$.
Next, we have calculated the $\pi^{0}\pi^{0}$ and $\pi^{0}\bar{K}^{0}$ invariant mass distribution with the fitted parameters in Fig. 4. The black points with error bars labeled as ‘CLEO data’ are the CLEO-c data taken from Ref. CLEO:2011cnt . One can find that, in the $\pi^{0}\pi^{0}$ invariant mass distribution, our results could well reproduce the structure of the cusp followed by a dip, which could be associated with the scalar meson $f_{0}(980)$, dynamically generated from the $S$-wave pseudoscalar meson pseudoscalar meson interactions. Meanwhile, our results show that the peak around $M^{2}_{\pi^{0}\pi^{0}}=0.2$ GeV2 is mainly due to the scalar meson $f_{0}(500)$ and the reflection of the vector meson $\bar{K}^{*0}$. Within the coupled-channel Bethe-Salpeter approach with the meson-meson potentials provided by the lowest order chiral lagrangian Oller:1997ti , one can simultaneously find a broad resonance $f_{0}(500)$, which couples strongly to $\pi\pi$ channel, and the one $f_{0}(980)$, which couples strongly to $K\bar{K}$ channel, which implies that these two states have a different dynamically internal structure.
In addition, our results for the $\pi\bar{K}^{0}$ invariant mass distribution agree well with the CLEO measurements. The narrow peak structure around $M^{2}_{\pi^{0}\bar{K}^{0}}=0.7$ GeV2 should be associated with the vector meson $K^{*}{(892)}$, while the contribution from the $S$-wave pseudoscalar meson-pseudoscalar meson interaction serves as a smooth background in the $\pi\bar{K}^{0}$ invariant mass distribution.
@@
We also present the Dalitz plot of ‘${M^{2}_{\pi^{0}\bar{K}^{0}}}$’ and ‘${M^{2}_{\pi^{0}\pi^{0}}}$’ in Fig. 5, which clearly shows the structures of $f_{0}(980)$ and $\bar{K}^{*}(892)$. One can also find that the peak near the $\pi^{0}\pi^{0}$ threshold is mainly due to the scalar meson $f_{0}(500)$ and the reflection of the vector meson $\bar{K}^{*0}$.
Finally, with the fitted parameters, we could estimate that the $S$-wave $\pi^{0}\pi^{0}$ interaction contributes about $47.7\%$ of the total decay rate, larger than the experimental data $(28.9\pm 6.3\pm 3.1)\%$, while the intermediate resonance $K^{*}(892)$ contributes about $52.3\%$ of the total decay rate, consistent with the experimental data $(65.6\pm 5.3\pm 2.5)\%$ within the uncertainties CLEO:2011cnt . It should be pointed out that the contribution from the scalar $f_{0}(500)$ was not taken into account by CLEO CLEO:2011cnt , which plays an important role in the near-threshold enhancement of the $\pi^{0}\pi^{0}$ invariant mass distribution.
Figure 5: The Dalitz plot of ’$M^{2}_{\pi^{0}\pi^{0}}$’ vs.’$M^{2}_{\pi^{0}\bar{K}^{0}}$’ for the process $D^{0}\to\pi^{0}\pi^{0}\bar{K}^{0}$.
@@ IV SUMMARY AND CONCLUSION
Motivated by the CLEO measurements on the process $D^{0}\to\pi^{0}\pi^{0}\bar{K}^{0}$ CLEO:2011cnt , we have investigated this process by taking into account the contribution from the $S$-wave pseudoscalar meson-pseudoscalar meson interaction within the chiral unitary method, which could dynamically generate intermediate resonance states $f_{0}(980)$ and $f_{0}(500)$, and also the contribution from the intermediate vector meson $\bar{K}^{*}{(892)}$.
First, we have performed a fit to the experimental data of the $\pi^{0}\pi^{0}$ and $\pi^{0}K_{S}^{0}$ invariant mass distributions. With the fitted parameters, our results could well reproduce the cusp structure followed a dip around 1 GeV in the $\pi^{0}\pi^{0}$ invariant mass distribution, which is mainly due to the contribution of scalar $f_{0}(980)$. Our results indicate that the peak near the $\pi^{0}\pi^{0}$ threshold is due to the scalar meson $f_{0}(500)$ and the reflection of the $\bar{K}^{*}$. Our results of the $\pi^{0}\bar{K}^{0}$ invariant mass distribution are in good agreement with the CLEO data, which supports that the peak around 892 MeV in the $\pi^{0}\bar{K}^{0}$ invariant mass distribution should be associated with $\bar{K}^{*}{(892)}$.
As we discussed in the introduction, there have different explanations for the nature of the scalar mesons $f_{0}(500)$ and $f_{0}(980)$. Our work, taking into account $S$-wave pseudoscalar meson-pseudoscalar meson interaction within the chiral unitary method, could well describe the CLEO measurements of the process $D^{0}\to\pi^{0}\pi^{0}\bar{K}^{0}$, especially the cusp structure followed a dip around 1 GeV in the $\pi^{0}\pi^{0}$ invariant mass distribution, which favors the molecular explanation of the scalar mesons $f_{0}(500)$ and $f_{0}(980)$.